OpenStudy (anonymous):

PLEASE HELP Evaluate the expression: v ⋅ w Given the vectors: r = <5, -5, -2>; v = <2, -8, -8>; w = <-2, 6, -5>

imqwerty (imqwerty):

-12 :)

OpenStudy (anonymous):

how did you get that? @imqwerty

imqwerty (imqwerty):

we need to find v.w we have- v=2i-8j-8k j=-2i+6j-5k for any 2 vectors a,b a.b= (product of their component along i)+(product of their component along j)+(product of their component along k) so v.j=(2x -2)+(-8x6)+(-8x -5) =-12

OpenStudy (anonymous):

ohh okay thank you so much, could you help me with one other question

imqwerty (imqwerty):

ok

OpenStudy (anonymous):

Find a ⋅ b. a = <7, 4>, b = <3, 5>

imqwerty (imqwerty):

have u tried solving this problem??

imqwerty (imqwerty):

41

OpenStudy (anonymous):

i got 21,20

imqwerty (imqwerty):

well a=7i+4j b=3i+5j a.b=(7x3)+(4x5) =21+20 =41